OpenStudy (anonymous):

What is the recursive formula for the sequence? 15,26,48,92

OpenStudy (jim766):

set up a table x y 1 15 2 26 3 48 4 92 a formula that would work would be y = 11x + 4

OpenStudy (jim766):

15 = 11(1) + 4 26 = 11(2) + 4 37 = 11(3) + 4 I hope you forgot the 37 : ) 48 = 11(4) + 4

OpenStudy (jim766):

recursive an = an-1 - 11 a2 = 15 + 11 a3 = 26 + 11